# Needle: A Generative-AI Powered Monte Carlo Method for Answering Complex Natural Language Queries on Multi-modal Data

### v1

@@ Abstract.
Multi-modal data, such as image data sets, often miss the detailed descriptions that properly capture the rich information encoded in them. This makes answering complex natural language queries a major challenge in these domains. In particular, unlike the traditional nearest-neighbor search, where the tuples and the query are modeled as points in a data cube, the query and the tuples are of different natures, making the traditional query answering solutions not directly applicable for such settings. Existing literature addresses this challenge for image data through vector representations jointly trained on natural language and images. This technique, however, underperforms for complex queries due to various reasons.
@@
This paper takes a step towards addressing this challenge by introducing a Generative-AI (GenAI) powered Monte Carlo method that utilizes foundation models to generate synthetic samples that capture the complexity of the natural language query and transform it to the same space of the multi-modal data. Following this method, we develop a system for image data retrieval and propose practical solutions that enable leveraging future advancements in GenAI and vector representations for improving our system’s performance. Our comprehensive experiments on various benchmark datasets verify that our system significantly outperforms state-of-the-art techniques.
Image Databases; Generative AI; Foundation Models;
1. Introduction
a Needle in a haystack: something that is impossible or extremely difficult to find, especially because the area you have to search is too large.”
Cambridge Dictionary
Figure 1. The query results for the queries: (a) [a banana] (row 1), (b) [an unripe banana] (row 2), and (c) [a banana gazing at its reflection in a mirror] (row 3), using Needle (first column) v.s. CLIP (second column).
@@ 1. Introduction
Multi-modal data, including image datasets, pose new challenges for data management systems. Unlike traditional databases, where tuple attributes are explicitly specified, multi-modal data often miss proper descriptions that capture the rich information encoded in them. As a result, traditional query-answering approaches are not directly applicable in these settings. On the other hand, advanced personal devices such as cell phones have enabled the collection of large volumes of multi-modal data, even at the individual user level. Due to the wealth of information “hidden” in each multi-modal tuple, the users may compose complex natural language queries, describing the tuple(s) they are interested in. To further motivate this, let us consider the following running example.
Example 1.1.
@@
A photo enthusiast has collected a large pool of images over the past several years on her private data repository. She wants to retrieve a specific photo she has in mind. She describes the picture as [A banana gazing at its reflection in a mirror].
@@
Traditional nearest-neighbor (NN-) search algorithms are not directly applicable here since those consider the database tuples and the query as points in a data cube. In contrast, in Example 1.1, (i) each tuple is an image, and (ii) the query is in the form of an unstructured, complex natural language statement.
@@
The state-of-the-art approach for addressing the need for answering natural language queries on image datasets is CLIP (Radford et al., 2021), a vector representation (aka embedding) jointly trained on several hundred million (image, text) pairs. Using CLIP as the embedder, one can transform the query and the dataset images to vector representation in the same space and apply NN-search to answer the query. This approach works satisfactorily for “simple” queries, such as simple object detection tasks. To see a concrete example, following with our running example (Example 1.1), we used a benchmark image dataset (LVIS (Gupta et al., 2019)) and issued the query [A banana]. As reflected in the top-right (row 1, col. 2) cell of Figure 1, the results were satisfactory, as most of the returned results include a banana (it also returned some irrelevant results, such as an image of a person wearing a banana costume). However, it failed for a moderately complex query, [An unripe banana], as none of the returned images was an unripe (green) banana – the middle-right (row 2, col. 2) cell of Figure 1. It further failed for the more complex queries, including the one in Example 1.1, and was not able to find the image of interest in its results. – the bottom-right (row 3, col. 2) cell of Figure 1. As we shall further explain in § 6, we observed a consistent behavior for this approach across a wide range of queries.
@@
This motivates us to take a step toward filling this gap in this paper by proposing a Monte Carlo randomized algorithm. In particular, we empower our algorithm with generative AI (GenAI) to generate a set of synthetic multi-modal tuples that represent the complex natural language query. Then, leveraging a collection of embedders, we conduct a series of NN-search using the generated tuples and aggregate the results to generate the final query outputs. We use our Monte Carlo method to develop Needle, our system for complex natural query answering on image datasets.
@@
Unlike existing approaches, Needle is capable of handling natural language queries with different levels of complexity. To demonstrate this, let us consider our running example, Example 1.1 once again. To compare Needle with CLIP, we issued the same set of simple, moderate, and complex queries. The results are provided in the first column of Figure 1. For the simple object-finding task, the top-left (row 1, col 1) cell in Figure 1, we observed a similar (but slightly better) performance since all returned images satisfied the query, i.e., they included a banana. It also performed reasonably well for the moderately complex query, as a large portion of the output images were green bananas – the middle-left (row 2, col 1) cell in Figure 1. But perhaps, more interestingly, it could successfully find the corresponding image described in the complex query of Example 1.1 – the bottom-left (row 3, col 1) cell in Figure 1. Interestingly, the target image was returned as the top-1 search result (the top-left image in the bottom-left cell), demonstrating the ability of our system to answer these types of queries. We observed a consistent behavior for our system in our experiments across a large number of queries (§ 6).
Summary of contributions
@@
This paper introduces a new approach to using GenAI for complex natural-language query answering on multi-modal data. Accordingly, it proposes Needle, a system for image data retrieval. In summary, our contributions are the following:
@@
  • We introduce a novel Monte Carlo method for answering complex natural-language queries on multi-modal data (§ 3). We leverage Foundation models to generate synthetic tuples representing natural language queries. We then use a collection of vector representations (aka embeddings) that allows us to apply traditional k-NN techniques to retrieve related tuples. Our Monte Carlo algorithm aggregates the output generated for each (synthetic tuple, embedding) pair to obtain the final result.
  • We use our Monte Carlo method to develop Needle, our interactive system image data retrieval (§ 4).
  • We propose a tiling technique, Smart Tiling, inspired by quadtrees (Finkel and Bentley, 1974), to detect and index objects in a densely populated image during the preprocessing time and tile the image based on the density of each image region (§ 4.2).
  • We propose a method to aggregate the top-$k$ rankings generated by each of the (synthetic tuple, embedding) pairs into a single final ranking while accounting for the reliability of each embedder (§ 5.1). This method enables readily adding/removing any of the current and future embedders to Needle.
  • We propose a LOF-based (Breunig et al., 2000) approach for detecting the poorly generated synthetic tuples as anomalies and discarding them (§ 5.2).
  • We conduct comprehensive experiments on multiple datasets to demonstrate the effectiveness of Needle in object detection benchmarks (§ 6.2.1), and its ability to handle scenarios with targeted queries (§ 6.2.2). We also explore the effect of different hyperparameters on Needle’s performance (§ 6.3) and the effectiveness of built-in anomaly detection method (§ 6.4).
  • Lastly, we conduct a case study where human evaluators compare Needle and OpenAI’s CLIP results, with 70% of participants preferring Needle over CLIP for their queries (§ 6.5).
2. Preliminaries
@@
Data Model: We consider the dataset $\mathcal{D}=\{t_{1},t_{2},\ldots,t_{n}\}$ as a set of $n$ multi-modal tuples (e.g., images) with no explicit values on specific attributes or any meta-data description about them. This model complies with real-world needs, where personal devices such as cell phones have advanced cameras, enabling a vast volume of multi-modal data collection and sharing by any user of such devices.
@@
Query Model: The input query $\varphi$ is defined as a natural language description that corresponds to a subset of tuples $\mathcal{D}_{\varphi}\subseteq\mathcal{D}$, where $\mathcal{D}_{\varphi}=\{t^{\prime}_{1},t^{\prime}_{2},\ldots,t^{\prime}_{k}\}$ represents the target subset intended for retrieval.
3. The GenAI-powered Monte-Carlo Method
Nearest-neighbor search in traditional settings is a well-studied line of research where many algorithms and indices such as Voronoi diagrams (Kolahdouzan and Shahabi, 2004), tree data structures (Yianilos, 1993), and local sensitive hashing (Jafari et al., 2021; Huang et al., 2015) have been proposed for efficient exact and approximate query answering. In such settings, the data points are specified as points in a $d$ dimensional space, i.e., $t_{i}\in\mathbb{R}^{d}$, where the goal is to find the $k$ nearest neighbors to a query point $q\in\mathbb{R}^{d}$ based on a distance function $dist:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$. That is, $k\mbox{-min}_{t_{i}\in\mathcal{D}}dist(q,t_{i})$. Such approaches, however, are not directly applicable in our setting because of the following challenges.
  1. (1)
    Multi-modal tuples such as images are not represented as points in $\mathbb{R}^{d}$.
  2. (2)
    The query is in the form of a (possibly complex) natural language phrase.
Our key idea for resolving these challenges is to leverage recent advancements in Generative AI (GenAI) and develop a randomized Monte Carlo algorithm that enables applying traditional nearest-neighbor techniques in our setting.
3.1. Vector Representations
@@
Addressing the first challenge is possible using the vector representation (aka embedding) of the tuples (Oquab et al., 2014; Kiela and Bottou, 2014). A vector representation $\mathcal{E}:dom(t)\rightarrow\mathbb{R}^{d_{\mathcal{E}}}$ is a transformation of the tuples to a high-dimensional vector of numeric values where the semantic similarity of tuples is proportional to the cosine similarity of their vector representations. Specifically, given a vector representation $\mathcal{E}^{\ell}$, we present $\mathcal{E}^{\ell}(t_{i})$ as $\vec{v}^{\ell}_{i}=\langle v_{1},v_{2},\cdots,v_{d_{\mathcal{E}^{\ell}}}\rangle$. The distance between two vector embeddings $\vec{v}^{\ell}_{i}$ and $\vec{v}^{\ell}_{j}$ is computed based on their cosine similarity as
$\delta^{\ell}_{i,j}=\delta(\vec{v}^{\ell}_{i},\vec{v}^{\ell}_{j})=1-\cos\left(\angle(\vec{v}^{\ell}_{i},\vec{v}^{\ell}_{j})\right)$
Let $\Delta(t_{i},t_{j})$ be the (unknown) semantic distance of the tuples $t_{i}$ and $t_{j}$.
@@
A vector representation $\mathcal{E}^{o}$ is optimal if it captures the semantic similarity of tuples, i.e., $\delta^{o}_{i,j}\sim\Delta(t_{i},t_{j}),~{}\forall t_{i},t_{j}$. Formally,
(1) $\Delta(t_{i},t_{j})\geq\Delta(t_{i},t_{k})~{}\Leftrightarrow~{}\delta^{o}_{i,j}\geq\delta^{o}_{i,k}$
@@
3.2. A Baseline Approach
The state-of-the-art resolution for the second challenge is to develop vector representations that are jointly trained on the multi-modal data and textual data. Specifically, CLIP (Radford et al., 2021) is an embedding trained jointly on (images, text) pairs. As a result, CLIP transforms text and image to the vector-representation same space. Subsequently, given a natural language query, one can compute its CLIP embedding and find its nearest neighbor images in the embedding space. This approach, however, suffers from two major issues:
First, the paired text with the images are usually simple sentences (e.g., [A photo of a dog]), missing to fully describe the rich information encoded in the image. Conversely, our goal is to answer complex natural language queries, describing a scene with multiple objects and their relationships (e.g., [A dog sitting by the fireplace, staring at the cat on the couch]). As a result, as we shall verify it in our experiments (§ 6), while such embeddings may perform reasonably on simple natural language queries their performance drops significantly on complex queries.
Second, as explained in Section 3.1, learned embeddings only preserve semantic similarities in expectation. As a result, relying solely on one vector representation for query answering may be misleading and inaccurate. Therefore, we instead develop a randomized Monte Carlo algorithm that utilized a set of vector representation transformations to develop more accurate query answering.
3.3. Query Transformation using GenAI
In traditional nearest neighbor problems both the dataset tuples and the query are represented as points in the same space. However, in our problem the query is form of a natural language phrase, while the dataset contains multi-modal tuples.
Let $\varphi$ be the given natural language query, and let $\{\vec{v}^{o}_{1},\cdots,\vec{v}^{o}_{n}\}$ be the vector representations for the tuples in $\mathcal{D}$. If we could specify the corresponding vector $\vec{q}^{o}(\varphi)$ for $\varphi$ in the $\mathcal{E}^{o}$ space, we could find the $k$-nearest neighbors of $\varphi$ as
(2) $k\mbox{-NN}(\varphi,\mathcal{D})=\underset{t_{i}\in\mathcal{D}}{k\mbox{-argmin}}~{}\delta\left(\vec{q}^{o}(\varphi),\vec{v}^{o}_{i}\right)$

### v2

# Needle: A Generative AI-Powered Multi-modal Database for Answering Complex Natural Language Queries

@@ Abstract
Multi-modal datasets, like those involving images, often miss the detailed descriptions that properly capture the rich information encoded in each item. This makes answering complex natural language queries a major challenge in this domain. In particular, unlike the traditional nearest neighbor search, where the tuples and the query are represented as points in a single metric space, these settings involve queries and tuples embedded in fundamentally different spaces, making the traditional query answering methods inapplicable. Existing literature addresses this challenge for image datasets through vector representations jointly trained on natural language and images. This technique, however, underperforms for complex queries due to various reasons.
@@
This paper takes a step towards addressing this challenge by introducing a Generative-based Monte Carlo method that utilizes foundation models to generate synthetic samples that capture the complexity of the natural language query and represent it in the same metric space as the multi-modal data.
Following this method, we propose Needle, a database for image data retrieval. Instead of relying on contrastive learning or metadata-searching approaches, our system is based on synthetic data generation to capture the complexities of natural language queries. Our system is open-source and ready for deployment, designed to be easily adopted by researchers and developers. The comprehensive experiments on various benchmark datasets verify that this system significantly outperforms state-of-the-art text-to-image retrieval methods in the literature. Any foundation model and embedder can be easily integrated into Needle to improve the performance, piggybacking on the advancements in these technologies.
1 Introduction
@@
Multi-modal datasets, like images, pose new challenges for data management systems. Unlike traditional databases, where tuple attributes are explicitly specified, these datasets often miss proper descriptions that capture the rich information encoded in each item. As a result, traditional query-answering approaches are not directly applicable in these settings. Meanwhile, modern personal devices like smartphones have made it possible to collect large amounts of multi-modal data, even from individual users. Due to the wealth of information “hidden” in each multi-modal tuple, the users may compose complex natural language queries, searching for the tuple(s) they are interested in. To further motivate this, let us consider the following running example.
@@
A photo enthusiast has collected a large pool of images over the past several years in her private data repository. She wants to retrieve a specific photo she has in mind. She describes the picture as [A banana gazing at its reflection in a mirror]. [^1] See Figure 3 for more examples.
Figure 1: The query results for (a) [a banana] (row 1), (b) [an unripe banana] (row 2), (c) [a banana gazing at its reflection in a mirror] (row 3), using Needle vs. CLIP [53].
@@
Traditional Nearest Neighbor (NN) search algorithms are not directly applicable here since they consider the dataset tuples and the query as points in the same data space. In contrast, in Example 1, (i) each tuple is an image, and (ii) the query is in the form of an unstructured, complex natural language statement.
[…]
@@
As a concrete example, continuing with our running example (Example 1), we use CLIP [53], OpenAI’s well-known jointly-trained embedding model, to answer the natural language query [A banana] on the benchmark image dataset LVIS [22]. As reflected in the top-right (row 1, col. 2) cell of Figure 1, the results are satisfactory, as most of the returned images include a banana. However, it fails for a moderately complex query, [An unripe banana], as none of the returned images shows an unripe (green) banana – the middle-right (row 2, col. 2) cell of Figure 1. It also fails on more complex queries, such as the one in Example 1, and does not retrieve the target image, the one shown in the bottom-right cell (row 3, col. 2) of Figure 1. As we further explain in § 5, we observe a similar behavior to this example across a wide range of queries.
@@
This motivates us to take a step toward filling this gap by proposing a Monte Carlo randomized algorithm. In particular, we empower our algorithm with generative AI (GenAI) to generate a set of synthetic multi-modal tuples that represent the complex natural language query. Subsequently, leveraging a collection of embedders, we conduct a series of NN searches using the generated tuples and aggregate the results to generate the final query outputs. We use our method to develop Needle  [^2] The name Needle is inspired by the common expression ”finding Needle in the haystack,” reflecting the challenge of pinpointing relevant information in large multimodal spaces., our deployment-ready open-source system333 Needle Github: https://github.com/UIC-InDeXLab/Needle/tree/main/ for complex natural query answering on image datasets.
@@
Unlike existing approaches, Needle can handle natural language queries with different levels of complexity. To demonstrate this, let us consider our running example, Example 1 once again. We issued the same set of simple, moderate, and complex queries to Needle. The results are provided in the first column of Figure 1. For the simple object-finding task, the top-left (row 1, col 1) cell in Figure 1, we observed a similar (but slightly better) performance since all returned images satisfied the query, i.e., they included a banana. It also performed reasonably well for the moderately complex query, as a large portion of the output images were green bananas – the middle-left (row 2, col 1) cell in Figure 1. But perhaps, more interestingly, it could successfully find the image-of-interest described in the complex query of Example 1 – the bottom-left (row 3, col 1) cell in Figure 1. Interestingly, the target image was returned as the top-1 search result (the top-left image in the bottom-left cell), demonstrating the ability of our system to answer these types of queries. We observed a consistent behavior for our system in experiments across a large number of different queries (§ 5).
@@
Summary of contributions: This paper introduces a new approach to using GenAI for complex natural-language query answering on multi-modal data. We propose Needle, a system for image data retrieval. In summary, our contributions are the following:
@@
  • We introduce a novel Monte Carlo method for answering complex natural-language queries on multi-modal data. We leverage Foundation models to generate synthetic tuples representing the queries. We then use a collection of vector representations (aka embeddings) that allows us to apply traditional k-NN techniques to retrieve related tuples. Our Monte Carlo algorithm aggregates the output generated for each (synthetic tuple, embedding) pair to obtain the final result. (§ 3)
  • We present Needle, a concrete, interactive system built upon our Monte Carlo method. Developed with production readiness in mind, Needle emphasizes ease of deployment, efficiency, robustness, and user-friendly interaction via a command-line interface (§ 4, Appendix D).
  • We propose and integrate several practical optimizations designed to significantly improve Needle’s performance and efficiency, including a dynamic embedder trust mechanism, outlier detection for generated data, caching, and a query complexity classifier (§ 3.3).
  • We conduct a comprehensive experimental evaluation demonstrating Needle’s effectiveness across diverse benchmarks (including object detection and complex natural language queries) compared to state-of-the-art baselines, supported by hyperparameter analysis and a human study indicating strong user preference for Needle (§ 5.2.1, § 5.2.2, Appendix E.4, Appendix E.5).
2 Related Work
Our paper relates to the following literature:
Multi-modal Data Retrieval: Multi-modal data retrieval (aka cross-modal retrieval), integrating various data types, has recently attracted significant attention. Early methods focused on shared representations by mapping modalities into a common latent space using canonical correlation analysis (CCA) [55, 70] and kernel-based techniques [26]. Recent approaches leverage deep learning models, including cross-modal encoders [69, 39] and joint embedding networks [6, 58]. Advanced techniques, such as attention mechanisms [76, 62] and contrastive learning [52], further enhance retrieval by capturing fine-grained relationships. CLIP [53] and ALIGN [25] are widely adopted methods for cross-modal retrieval, both of which are based on contrastive learning approaches. Feedback mechanisms [48, 84] and user preferences [81, 11] improve multi-modal search results. In addition, there exists practical approaches to enrich the semantic metadata of images using image tagging [20, 83, 35] and object detection techniques [85]. Research also extends to other modalities, such as text-video retrieval [36, 43, 64, 47, 46, 40] and text-audio retrieval [74, 10, 28]. For a detailed survey of cross-modal retrieval methods and future directions, see [33, 71].
To the best of our knowledge, this paper is the first to introduce the idea of utilizing GenAI to synthetically generate samples that capture natural language complexities and proposes a GenAI-powered Monte Carlo method for multi-modal data retrieval.
GenAI in Data Management: The intersection of Generative AI (GenAI) and data management has emerged as a promising area, driving innovations in how data is generated, managed, and utilized. GenAI models, such as GPT [1] and Stable Diffusion [57], have been employed for synthetic data generation to augment datasets [61, 59], improve model training [30, 65], and address data scarcity or imbalance issues [16]. In data cleaning and integration, these models assist with tasks like imputation [68], and anomaly detection by generating plausible data patterns [77]. Recent works also explore the role of GenAI in automating query generation [66, 67, 37] and data summarization [14], enhancing the efficiency of database systems and analytics platforms.
Natural Language Query Processing: The evolution of natural language to structured query translation has advanced from rule-based systems to sophisticated large language models (LLMs). Early methods used template matching and predefined rules [3, 32], which were inflexible and required frequent manual updates. Machine learning approaches introduced more adaptability by learning query patterns from annotated datasets [41, 82], though they struggled with complex language constructs. Deep learning models, especially transformers like BERT [15] and GPT [9], enabled semantic understanding and laid the foundation for text-to-SQL systems [34]. Modern LLMs, such as PaLM [12] and GPT-4 [1], generate context-aware queries across multiple domains with minimal supervision. Retrieval-augmented generation (RAG) further enhances query precision by leveraging external knowledge [31]. Ongoing efforts focus on fine-tuning LLMs to improve accuracy and efficiency in converting natural language to structured queries [66, 67, 37].
3 The GenAI-powered Monte-Carlo Method
@@
We consider the dataset $\mathcal{D}=\{t_{1},t_{2},\ldots,t_{n}\}$ as a set of $n$ multi-modal tuples with no explicit values on specific attributes or meta-data description about them. This model complies with real-world needs, where personal devices such as cell phones have advanced cameras, enabling a vast volume of multi-modal data collection and sharing by any user of such devices.
@@
The input query $\varphi$ is defined as a natural language description that corresponds to a subset of tuples $\mathcal{D}_{\varphi}\subseteq\mathcal{D}$, where $\mathcal{D}_{\varphi}=\{t^{\prime}_{1},\cdots,t^{\prime}_{k}\}$ represents the target subset for retrieval.
[…]
@@
Vector Representations: A vector representation $\mathcal{E}:dom(t)\rightarrow\mathbb{R}^{d_{\mathcal{E}}}$ is a transformation of the tuples to a high-dimensional vector of numeric values where the semantic similarity of tuples is proportional to the cosine similarity of their vector representations. Specifically, given a vector representation (aka embedder) $\mathcal{E}^{\ell}$, we present $\mathcal{E}^{\ell}(t_{i})$ as $\vec{v}^{\ell}_{i}=\langle v_{1},v_{2},\cdots,v_{d_{\mathcal{E}^{\ell}}}\rangle$. The distance between two vector embeddings $\vec{v}^{\ell}_{i}$ and $\vec{v}^{\ell}_{j}$ is computed based on their cosine similarity as
$\delta^{\ell}_{i,j}=\delta(\vec{v}^{\ell}_{i},\vec{v}^{\ell}_{j})=1-\cos\left(\angle(\vec{v}^{\ell}_{i},\vec{v}^{\ell}_{j})\right)$ (1)
@@
Let $\Delta(t_{i},t_{j})$ be the (unknown) semantic distance of the tuples $t_{i}$ and $t_{j}$. A vector representation $\mathcal{E}^{o}$ is optimal if it captures the semantic similarity of tuples, i.e., $\delta^{o}_{i,j}\sim\Delta(t_{i},t_{j}),~{}\forall t_{i},t_{j}$. Formally,
$\Delta(t_{i},t_{j})\geq\Delta(t_{i},t_{k})~{}\Leftrightarrow~{}\delta^{o}_{i,j}\geq\delta^{o}_{i,k}$ (2)
@@
A Baseline Approach: A vector representation that is jointly trained on the multi-modal data and textual data transform text and image to the vector-representation space. Therefore, given a natural language query, one can compute its embedding and find its nearest neighbor images in the embedding space. This approach, however, suffers from two major issues:
First, the paired texts with the images are usually simple sentences (e.g., [A photo of a dog]), unable to fully describe the rich information encoded in the image. Conversely, our goal is to answer complex natural language queries, describing a scene with multiple objects and their relationships (e.g., [A dog sitting by the fireplace, staring at the cat on the couch]). As a result, as we shall verify in our experiments (§ 5), while such embeddings may perform reasonably on simple natural language queries, their performance drops significantly on complex queries.
Second, as explained earlier, learned embeddings only preserve semantic similarities in expectation. As a result, relying solely on one vector representation for query answering may be misleading and inaccurate. Hence, we instead utilize an ensemble (a set) of vector representation transformations, possibly trained on different training datasets, to develop our Monte Carlo algorithm for more accurate query answering.
3.1 Query Transformation using GenAI
In traditional nearest-neighbor problems, both the dataset tuples and the query are represented as points in the same space. However, in our problem, the query is in the form of a natural language phrase, while the dataset contains multi-modal tuples.
Let $\varphi$ be the given natural language query, and let $\{\vec{v}^{o}_{1},\cdots,\vec{v}^{o}_{n}\}$ be the vector representations for the tuples in $\mathcal{D}$. If we could specify the corresponding vector representation $\vec{q}^{o}(\varphi)$ for $\varphi$ in the $\mathcal{E}^{o}$ space, we could find the $k$-nearest neighbors of $\varphi$ as
$\hskip 170.71652ptk\mbox{-NN}(\varphi,\mathcal{D})=\underset{t_{i}\in\mathcal{D}}{k\mbox{-argmin}}~{}\delta\left(\vec{q}^{o}(\varphi),\vec{v}^{o}_{i}\right)$ (3)